Find the bandwidth for the following situations if we need to modulate a 5-KHz voice.

a. AM

b. PM (setß=3)

c. PM (set p=1)


1
Expert's answer
2021-04-01T11:09:00-0400

Given the frequency of the voice signal -

"(f)=5kHz"

a) Bandwidth of the amplitude modulation "B_{AM}=2B"

"=2\times 5kHz"

"=10kHz"

b) Bandwidth required for the phase modulation

"B_{PM}=2(1+\beta)B"

Now, substituting the values,

="2(1+3)\times 5 kHz"

"=40kHz"

c) Bandwidth required for the phase modulation,

"B_{PM}=2(1+\beta)B"

Now, substituting the values,

"=2(1+1)\times 5kHz"

"=20kHz"